Fritz Co. uses the perpetual inventory system. The company purchased $19,000 of merchandise from Baker Co. Fritz also paid $550 freight to obtain the goods under terms FOB shipping point. All of the merchandise purchased was sold for $25,450 cash. The amount of gross margin for this merchandise is:  

A.

$6,000

B.

$5,900

C.

$6,450

D. $7.000
Answer #1

Cost of goods sold = Purchase+Freight in = 19000+550 = 19550

Gross margin = Sales-Cost of goods sold = 25450-19550 = 5900

So answer is b) $5900